Solve the following equation:
-x/6=-2/7

Found 2 solutions by rfer, algebrahouse.com:
Answer by rfer(16322)

-x=-2/7*6
-x=-12/7
x=1 5/7

Answer by algebrahouse.com(1659)

Solve the following equation:
-x/6 = -2/7

This is a proportion. To solve, just cross-multiply.

-x/6 = -2/7
-7x = 6(-2) {cross-multiplied}
-7x = -12 {multiplied 6 by -2}
x = 12/7 {divided both sides by -7}
x = 1 5/7 {wrote as a mixed number}
www.algebrahouse.com